OpenStudy (anonymous):

M is the midpoint of AB . The coordinates of A are (-2,3) and the coordinates of M are (1,0). Find the coordinates of B.

OpenStudy (ivettef365):

use midpoint formula

OpenStudy (anonymous):

I got (-1/2,3/2)

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

x1+x2 ------ 2

OpenStudy (ivettef365):

-2+1 -1 ----- = ---- 2 2

OpenStudy (ivettef365):

y1+y2 3+0 ------ = ==== = 3/2 2 2

OpenStudy (anonymous):

[1] (-1/2, 3/2) [2] (4,-3) [3] (-4,3) [4] (-5,6) Those are the choices

OpenStudy (anonymous):

I picked one but its wrong

OpenStudy (ivettef365):

oh, because points M (1,0) are the midpoints

OpenStudy (anonymous):

So how would I do iy?

OpenStudy (anonymous):

It*

OpenStudy (ivettef365):

use same formula x1+x2 -2 +x ------ = ------- = 1 x 2

OpenStudy (anonymous):

Why is the answer (4,-3)

OpenStudy (ivettef365):

so that is equal to -2+x = 2, so x = 2+2

OpenStudy (ivettef365):

now solve for y 3 +y ==== = 0 2

OpenStudy (ivettef365):

so you have 3+y=0, y=-3

OpenStudy (anonymous):

So where did the 4 come from?

OpenStudy (ivettef365):

you have -2+x ----- = 1 2

OpenStudy (anonymous):

I have no idea how you did any of this

OpenStudy (ivettef365):

first thing you do is multiply by 2 so you have (2) -2+x --- .----- = 1 . 2 1 2

OpenStudy (ivettef365):

then you have -2+x=2 +2 +2 ------------- x=4

OpenStudy (ivettef365):

I used the midpoint formula, which is x1 + x2 -------- 2